OpenStudy (anonymous):

x^2+x-6 factor as much as possible

jimthompson5910 (jim_thompson5910):

find two numbers that a) multiply to -6 (last term) AND b) add to 1 (the middle coefficient, note: x = 1x)

jimthompson5910 (jim_thompson5910):

what two numbers did you get?

OpenStudy (anonymous):

-2 and -3?

jimthompson5910 (jim_thompson5910):

close, but not quite

jimthompson5910 (jim_thompson5910):

-2 and -3 add to -2+(-3) = -5 -2 and -3 multiply to -2*(-3) = 6

jimthompson5910 (jim_thompson5910):

try again

jimthompson5910 (jim_thompson5910):

Hint: one of the numbers must be negative and the other must be positive (so they multiply to a negative number)

OpenStudy (anonymous):

2 and -3

jimthompson5910 (jim_thompson5910):

closer, but still not there yet

jimthompson5910 (jim_thompson5910):

2 + (-3) = -1 but we want positive 1

OpenStudy (anonymous):

i'm so sorry I know I sound really stupid right now but i'm horrible at this stuff

jimthompson5910 (jim_thompson5910):

its ok and you're not stupid

jimthompson5910 (jim_thompson5910):

that's what practice is for

jimthompson5910 (jim_thompson5910):

so get that silly idea out of your head, what else is left to try?

jimthompson5910 (jim_thompson5910):

another hint: since the two numbers add to a positive number, the larger number is positive (ignore the sign)

OpenStudy (anonymous):